Preble Company manufactures one product. Its variable manufacturing overhead is applied to production based on direct labor-hours and its standard cost card per unit is as follows:

Direct material: 6 pounds at $8.00 per pound $ 48.00
Direct labor: 4 hours at $13 per hour 52.00
Variable overhead: 4 hours at $5 per hour 20.00
Total standard variable cost per unit $ 120.00

The company also established the following cost formulas for its selling expenses:

Fixed Cost per Month Variable Cost per Unit Sold
Advertising $ 380,000
Sales salaries and commissions $ 460,000 $ 30.00
Shipping expenses $ 21.00

The planning budget for March was based on producing and selling 20,000 units. However, during March the company actually produced and sold 25,500 units and incurred the following costs:

  1. Purchased 170,000 pounds of raw materials at a cost of $7.20 per pound. All of this material was used in production.
  2. Direct-laborers worked 73,000 hours at a rate of $14.00 per hour.

  3. Total variable manufacturing overhead for the month was $427,050.

  4. Total advertising, sales salaries and commissions, and shipping expenses were $386,000, $545,000, and $295,000, respectively.

Required:

1. What raw materials cost would be included in the company’s flexible budget for March?

2. What is the materials quantity variance for March? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.). Input the amount as

3. What is the materials price variance for March? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.). Input the amount as a positive value.)

4. If Preble had purchased 188,000 pounds of materials at $7.20 per pound and used 170,000 pounds in production, what would be the materials quantity variance for March? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.). Input the amount as a positive value.)

5. If Preble had purchased 188,000 pounds of materials at $7.20 per pound and used 170,000 pounds in production, what would be the materials price variance for March? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.). Input the amount as a positive value.)
Answer #1
1
Materials cost in flexible budget 1224000 =25500*48
2
Materials quantity variance 136000 U =8*(170000-25500*6)
3
Materials price variance 136000 F =170000*(7.2-8)
4
Materials quantity variance 136000 U =8*(170000-25500*6)
5
Materials price variance 150400 F =188000*(7.2-8)